A car of mass 50tonnes pulls another car of mass 10tonnes, starting from rest the 2 masses attain a velocity of 20m/s in 5s. The coefficient of friction on each car equals 5N/tonne. Calculate

a) Force exerted by the engine

b) time required to cover 100m

 


1
Expert's answer
2021-08-04T04:40:30-0400

"\text{Let}"

"m_1 = 50 t = 50000kg"

"m_2 = 10 t = 100000kg"

"v_0=0;v_1=20\frac{m}{s};t=5s"

"\mu = 5\frac{N}{t}= 0.005\frac{N}{kg}"

"F_{ex}-\text{Force exerted by the engine}"

"F_{fr}-\text{Friction force}"

"F = ma;F = F_{ex}-F_{fr}"

"a = \frac{v_1-v_0}{t}= \frac{20-0}{5}=4\frac{m}{s^2}"

"F_{fr}= \mu*m"

"m = m_1+m_2"

"a)"

"F_{ex}=F+F_{fr}= ma+\mu m=200250N"

"b)"

"S = v_0t +\frac{at^2}{2}"

"v_0 = 0"

"t =\sqrt{\frac{2S}{a}}=\sqrt{\frac{2*100}{4}}=7.07\frac{m}{s}"


"\text{Answer:}"

"a) F_{ex}=200250N"

"b)t =7.07\frac{m}{s}"